OpenStudy (anonymous):

Aisha has 12 pages of math homework. If she takes 20 minutes to complete one page, how long will it take her to finish all 12 pages? Please help , I thought it was 1 hour, 20 mins. Am I correct ?

OpenStudy (anonymous):

if she takes 20 mins time for one page then for 2 pages she will take

OpenStudy (anonymous):

4 hours

OpenStudy (anonymous):

12 *20= 240 minutes or 4hours

OpenStudy (anonymous):

20 * 12 = 240 minutes = 240 / 60 = 4 hours

OpenStudy (anonymous):

12 *20= 240 minutes or 4hours

OpenStudy (anonymous):

i feel u dont want to understand u r only interested in answer

OpenStudy (anonymous):

12 *20= 240 minutes or 4hours

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

Oh Okay , Now I Understand. Thank You

OpenStudy (anonymous):

12 *20= 240 minutes or 4hours

OpenStudy (anonymous):

4 hrs

OpenStudy (anonymous):

Thank You (: